A uniform beam of length 7.60 m and weight 4.50 x 102 N is carried by two workers, Sam and Joe. Determine the force that each person exerts on the beam given that Sam is at a distance of 1.00 m away from one of the edge and Joe is 2.00 m towards the other edge of the beam. (Sam is walking towards Joe)
1
Expert's answer
2020-12-04T07:40:41-0500

We have given:

Weight=P=4.5"\times 102" N=459N,

Length=L=7.60m.

Let`s solve :

for Sam:

Fsam"\times" 4.60m=P"\times"1.8m,

Fsam=459"\times" 0.39=179.60N for Sam and

for Joe:

FJoe"\times" 4.60m=P"\times"2.8m,

FJoe=459"\times" 0.60=275.4N for Joe and

we have found answers.